### v1

# A Game-Theoretic Framework for Distributed Load Balancing: Static and Dynamic Game Models

Abstract
Motivated by applications in job scheduling, queuing networks, and load balancing in cyber-physical systems, we develop and analyze a game-theoretic framework to balance the load among servers in both static and dynamic settings. In these applications, jobs/tasks are often held by selfish entities that do not want to coordinate with each other, yet the goal is to balance the load among servers in a distributed manner. First, we provide a static game formulation in which each player holds a job with a certain processing requirement and wants to schedule it fractionally among a set of heterogeneous servers to minimize its average processing time. We show that this static game is a potential game and admits a pure Nash equilibrium (NE). In particular, the best-response dynamics converge to such an NE after $n$ iterations, where $n$ is the number of players. We then extend our results to a dynamic game setting, where jobs arrive and get processed in the system, and players observe the load (state) on the servers to decide how to schedule their jobs among the servers in order to minimize their averaged cumulative processing time. In this setting, we show that if the players update their strategies using dynamic best-response strategies, the system eventually becomes fully load-balanced and the players’ strategies converge to the pure NE of the static game. In particular, we show that the convergence time scales only polynomially with respect to the game parameters. Finally, we provide numerical results to evaluate the performance of our proposed algorithms under both static and dynamic settings.
{IEEEkeywords}
Games theory, Nash equilibrium, potential games, load balancing, job scheduling, distributed algorithms, best-response dynamics.
1 Introduction
Distributed and uncoordinated load balancing is challenging when a set of strategic agents compete for a set of shared resources. There is a large body of literature on load balancing, in which the proposed schemes can be either static or dynamic. In static load balancing schemes, agents use prior knowledge about the system, whereas dynamic load balancing methods make decisions based on the system’s real-time state.
Load balancing in strategic environments has been studied using game theory in the existing literature, where, in most applications, the system is often modeled as congestion games [1, 2]. In congestion games, each player’s utility is determined by the resources they select and the number of other players choosing the same resources. If all players choose the most beneficial resource, it becomes congested, and all players suffer from the congestion cost. As a result, agents may prefer to choose less profitable resources to avoid congestion. Moreover, uncoordinated competition over shared resources has also been analyzed through the lens of multi-armed bandit games [3, 4]. In such games, a set of independent decision-makers sequentially chooses among a group of arms with the goal of maximizing their expected gain, without having any prior knowledge about the properties of each arm.
One of the limitations of the aforementioned models is the assumption that subsequent rounds of these games are independent. In other words, the outcome of a given round has no effect on the outcome of future rounds [5]. However, there is a carryover effect between rounds in queuing systems, caused by the jobs remaining in the queues. In fact, the average wait time for each queue in each round depends not only on the number of tasks sent to it but also on the number of jobs remaining in the queue from previous rounds. Given these limitations and inspired by applications in job scheduling and queuing networks, we propose and analyze a game-theoretic framework to achieve load balancing between servers in both static and dynamic settings. In these scenarios, jobs are managed by self-interested players who are unwilling to coordinate; nonetheless, the objective is to distribute the load across servers in a balanced and decentralized manner.
We first introduce a static load balancing game model in which each player holds a job with a certain processing requirement and seeks to schedule it fractionally across a set of heterogeneous servers to minimize their average processing time. Next, to account for the carryover effect, we extend our model to a dynamic game setting, where jobs arrive and are processed in the system over time. In this setting, players observe the load (state) on the servers and decide how to schedule their jobs to minimize their cumulative average processing time. In both settings, we show that if players update their strategies using static or dynamic best-response at each state, the system eventually becomes load-balanced, and the players’ strategies converge to the pure Nash equilibrium (NE) of the underlying game. Notably, we show that the convergence time scales polynomially with respect to the game parameters in both settings, which is further justified using extensive numerical experiments.
1.1 Related Work
Load balancing in distributed systems has been extensively studied in the literature. Many of these studies focus on systems with a centralized scheduler [6, 7, 8, 9, 10]. Among the various load balancing approaches, join-the-shortest-queue (JSQ) stands out as one of the most thoroughly researched methods [11, 12, 13, 14, 15, 16, 17]. This approach assigns jobs to the server with the shortest queue length. Join-idle-queue (JIQ) [18, 19, 20] is another load balancing policy where jobs are sent to idle servers, if any, or to a randomly selected server. These methods might not perform well because the queue length is not always a good indicator of the servers’ wait times. Power-of-$d$-choices [9, 21, 22, 23, 24, 25, 26] queries $d$ randomly selected servers and sends the jobs to the least loaded ones. This method has been proven to yield short queues in homogeneous systems but may be unstable in heterogeneous systems.
Studying distributed load balancing through a game-theoretic lens is a natural approach. Despite its potential, relatively few studies have applied game theory to the load balancing problem [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37]. In [38], the authors model the load balancing problem in heterogeneous systems as a non-cooperative, repeated finite game and identify a structure for the best-response policy in pure strategies. They later introduce a distributed greedy algorithm to implement the best-response policy. However, since their setting is stateless, the proposed algorithm does not necessarily perform well in dynamic games. Moreover, [39] models the load balancing problem as a cooperative game among players.
The work [40] studies the distributed load balancing problem using multi-agent reinforcement learning methods. In this approach, each server simultaneously works on multiple jobs, and the efficiency with which the server handles a job depends on its capacity and the number of jobs it processes. The problem is modeled as a stochastic system where each player has access to local information. To evaluate server efficiency, each player maintains an efficiency estimator, a vector summarizing the server’s performance, which also serves as the player’s state. The players’ policies are then formulated as functions of these states. They compare the performance of their algorithm to other existing algorithms, such as non-adaptive algorithms—i.e., algorithms where players use servers in a predetermined manner—while revealing that naive communication between players does not necessarily improve the system’s overall efficiency.
Krishnasamy et al. [41] study a variant of the multi-armed bandit problem with the carryover effect, which is suitable for queuing applications. They consider a bandit setting in which jobs queue for service, and the service rates of the different queues are initially unknown. The authors examine a queuing system with a single queue and multiple servers, where both job arrivals to the queue and the service rates of the servers follow a Bernoulli distribution. At each time slot, only one server can serve the queue, and the objective is to determine the optimal server scheduling policy for each time slot. They further evaluate the performance of different scheduling policies against the no-regret policy [^1] The policy that schedules the optimal server in every time slot..
More recently, [5] studied the multi-agent version of the queuing system in [41]. Their setting differs from ours in the sense that, in our setting, each server has a separate queue and processes jobs with a fixed service rate. However, [5] describes a system where each player has a queue, and servers process jobs with a fixed time-independent probability. All unprocessed jobs are then sent back to their respective queues. Queues also receive bandit feedback on whether their job was cleared by their chosen server. One concern in such settings is whether the system remains stable, despite the selfish behavior of the learning players. This concern does not arise in settings with a centralized scheduler, as these systems remain stable if the sum of arrival rates is less than the sum of service rates. However, this condition does not guarantee stability in settings with strategic players. In [5], the authors show that if the capacity of the servers is high enough to allow a centralized coordinator to complete all jobs even with double the job arrival rate, and players use no-regret learning algorithms, then the system remains stable.A no-regret learning algorithm minimizes the expected difference in queue sizes relative to a strategy that knows the optimal server. They later show in [42] that stability can be guaranteed even with an extra capacity of $\frac{e}{e-1}\approx 1.58$, but only if players choose strategies that maximize their long-run success rate.
1.2 Contributions and Organization
This paper introduces a game-theoretic framework to address the problem of uncoordinated load balancing in systems with self-interested players. The key contributions of this work are as follows:
  • Static Game Model: We formulate a static game in which each player holds a job with specific processing requirements and aims to schedule it fractionally across a set of servers in order to minimize its average wait time. We show that this static game is a potential game, and hence admits a pure NE. Additionally, we show that the best-response dynamics converge to this pure NE after a number of iterations, where the number of iterations is equal to the number of players.
  • Dynamic Game Extension: To account for the carryover effect in queuing systems, we extend the static model to a dynamic game setting. In this extension, jobs arrive and are processed over time, and players update their scheduling strategies based on the current state of the servers. We show that if players use dynamic best-response strategies, the system eventually reaches a load-balanced state, and their strategies converge to the pure NE of the static game. Furthermore, we show that convergence occurs in a polynomial number of iterations.
  • Numerical Evaluation: We provide numerical results that evaluate the performance of our proposed algorithms in both static and dynamic settings, demonstrating the efficiency of our approach in achieving load balancing.
The remainder of the paper is organized as follows: In Section 2.1, we introduce the static game model, followed by its extension to a dynamic game setting in Section 2.2. We present our main results for the static game model in Section 3. In Section 4, we propose our algorithm for the dynamic game model and analyze its convergence properties. Section 5 presents numerical simulations to evaluate the performance of the proposed algorithms. Finally, Section 6 concludes the paper and discusses future research directions. Omitted proofs and other auxiliary results are deferred to the Appendix.
2 Problem Formulation
In this section, we first introduce the static load balancing game and subsequently extend it to a dynamic game setting.
2.1 Static Load Balancing Game
In the static load balancing game, there are a set $[m]=\{1,2,\ldots,m\}$ of $m$ servers with service rates $\mu_{j}>0$ $\forall j\in[m]$, and a set $[n]=\{1,2,\ldots,n\}$ of $n$ players. Each player $i\in[n]$ holds a job with processing requirement (length) $\lambda_{i}$. This job can be fractionally scheduled on different servers. [^2] Since there is a one-to-one correspondence between jobs and players, we often use the terms players and jobs interchangeably. We also define $\lambda_{\max}=\max_{i}\lambda_{i}$ and $\lambda_{\min}=\min_{i}\lambda_{i}$ to be the maximum and minimum length of jobs among all the players, respectively.
Let $a_{ij}$ be the fraction of job $i$ that is scheduled on server $j$ by player $i$. Then, the strategy (action) set for player $i$ is given by all the vectors $a_{i}=(a_{i1},\ldots,a_{im})$ that belong to the probability simplex
$A_{i}=\big{\{}a_{i}:\sum_{j=1}^{m}a_{ij}=1,a_{ij}\geq 0\big{\}}.$
We further assume that each server $j$ is initially occupied with a load of $s^{0}_{j}\geq 0$ jobs. A graphical illustration of the model is provided in Fig. 1. In this example, we consider three servers with initial loads $s_{1}^{0},s_{2}^{0},s_{3}^{0}$ and service rates $\mu_{1}=2$, $\mu_{2}=1.5$, and $\mu_{3}=1.0$, respectively. Two players are present, with job lengths $\lambda_{1}=3$ and $\lambda_{2}=5$. As shown in the figure, Player 1 has distributed 2 units of its job on Server 1, 0.5 units on Server 2, and the remaining 0.5 units on Server 3. Meanwhile, Player 2 schedules 1 unit of its job on Server 1 and the remainder on Server 3.
Now, let us suppose player $i$ has already scheduled $x\in[0,\lambda_{i})$ portion of its job on server $j$. The amount of time required by server $j$ to fully process that portion is given by $\frac{x+s^{0}_{j}}{\mu_{j}}$. In particular, if player $i$ schedules an additional infinitesimal $dx$ portion of its job on server $j$, the delay experienced by that infinitesimal portion to be processed is $\frac{x+s^{0}_{j}}{\mu_{j}}dx$. Therefore, the wait time for player $i$ when scheduling a fraction $\alpha\in[0,1]$ of its job on server $j$ can be calculated as:
[…]
Finally, since player $i$ schedules different fractions of its job on different servers according to its action $a_{i}$, the average wait time of player $i$ when taking action $a_{i}$ can be written as:
$\sum_{j=1}^{m}\lambda_{i}a_{ij}\left(\frac{\lambda_{i}a_{ij}}{2\mu_{j}}+\frac{s_{j}^{0}}{\mu_{j}}\right).$ (1)
Using the same reasoning, given an action profile $a=(a_{1},\ldots,a_{n})$ of all the players, we can define the cost function for player $i$ as
$D_{i}(a)=\sum_{j=1}^{m}\lambda_{i}a_{ij}\left(\frac{\lambda_{i}a_{ij}}{2\mu_{j}}+\frac{s^{0}_{j}+\sum_{k\neq i}\lambda_{k}a_{kj}}{\mu_{j}}\right).$ (2)
Therefore, we obtain a noncooperative static load balancing game $([n],\{D_{i}(a)\}_{i\in n},A_{1}\times\cdots\times A_{n})$, where each player $i$ aims to make a decision $a_{i}\in A_{i}$ in order to minimize its own cost function $D_{i}(a)$.
Figure 1: An illustration of the static load balancing game with two players and three servers.
2.2 Dynamic Load Balancing Game
The dynamic game setting is the same as the static one, except that the scheduling process evolves over time as more jobs arrive and get processed. More specifically, in the dynamic load balancing game, at each discrete time $t=0,1,2,\ldots$, we assume that only one player (say player $i$) receives a job [^3] For instance, if job arrivals follow a continuous random process such as a Poisson process, one may assume that at any time only one of the players receives a new job., and this player must immediately schedule it among the servers. Let us define $a_{i}^{t}$ and $s_{j}^{t}$ as the action of player $i$ and the load (state) of server $j$ at time $t$, respectively. Moreover, we denote the state of the game at time $t$ by $s^{t}=(s_{1}^{t},\ldots,s^{t}_{m})$, which is the vector of observed loads on each of the servers at time $t$. Now suppose that player $i$ receives a new job of length $\lambda_{i}$ at time $t$ and schedules it among the servers according to its action $a_{i}^{t}=(a_{i1}^{t},\cdots,a_{im}^{t})$. By following the same derivations as in the static game (Eq. (1)), the instantaneous average wait time for player $i$ due to its new job arrival at time $t$ equals
$D_{i}(a^{t},s^{t})=\sum_{j=1}^{m}\lambda_{i}a^{t}_{ij}\left(\frac{\lambda_{i}a^{t}_{ij}}{2\mu_{j}}+\frac{s^{t}_{j}}{\mu_{j}}\right).$ (3)
After that, the state of server $j\in[m]$ changes according to
$s_{j}^{t+1}=(s_{j}^{t}+\lambda_{i}a^{t}_{ij}-\mu_{j})^{+}\ \ \ \forall j\in[m],$ (4)
where $(\cdot)^{+}=\max\{0,\cdot\}$. Finally, the goal for each player $i$ is to follow a certain scheduling policy $\pi_{i}=(a^{t}_{i},t=1,2,\ldots)$ to minimize its overall averaged cumulative cost given by
$V_{i}(\pi_{i},\pi_{-i})=\limsup_{T\to\infty}\frac{1}{T}\sum_{t=1}^{T}D_{i}(a^{t},s^{t}).$ (5)
Assumption 1
In the dynamic load balancing game, we assume that $\lambda_{\max}<\sum_{j=1}^{m}\mu_{j}$.
We note that Assumption 1 is the stability condition and is crucial in the dynamic setting. Otherwise, one could always consider a worst-case instance where, regardless of any scheduling policy, the length of the waiting jobs in the server queues will grow unboundedly.
3 Main Theoretical Results for the Static Load Balancing Game
In this section, we analyze the static load balancing game and show that it belongs to the class of potential games, for which the best response dynamics are known to converge to a pure Nash equilibrium [43]. We then characterize the best response dynamics in a closed-form as a solution to a convex program. This closed-form solution will be used in the subsequent section to extend our results to the dynamic game setting.
Theorem 1
The static load balancing game is a potential game and admits a pure NE. In particular, the best response dynamics, in which players iteratively and sequentially redistribute their jobs over the servers by best responding to others’ strategies, converge to a pure NE.
Proof 3.2.
We show that the static load balancing game is an exact potential game [43], that is, we show that for any two action profiles $(a_{i},a_{-i})$ and $(a^{\prime}_{i},a_{-i})$ that only differ in the action of player $i$, we have
$\Phi(a_{i},a_{-i})-\Phi(a^{\prime}_{i},a_{-i})=D_{i}(a_{i},a_{-i})-D_{i}(a^{\prime}_{i},a_{-i}).$ (6)
To that end, let us consider the potential function
$\Phi(a)=\sum_{j=1}^{m}\frac{\mu_{j}}{2}\left(\frac{s^{0}_{j}+\sum_{k=1}^{n}\lambda_{k}a_{kj}}{\mu_{j}}\right)^{2}.$
For simplicity, let $s_{ij}=s^{0}_{j}+\sum_{k\neq i}\lambda_{k}a_{kj}$ be the effective load on server $j$ viewed from the perspective of player $i$, where we note that since $s_{ij},j\in[m]$ do not depend on player $i$’s action, they must be the same for both action profiles $(a_{i},a_{-i})$ and $(a^{\prime}_{i},a_{-i})$. We can write
$\Phi(a_{i},a_{-i})=$ $\sum_{j=1}^{m}\frac{\mu_{j}}{2}\left(\frac{s^{0}_{j}+\sum_{k=1}^{n}\lambda_{k}a_{kj}}{\mu_{j}}\right)^{2}$
$=$ $\sum_{j=1}^{m}\frac{\mu_{j}}{2}\left(\frac{\lambda_{i}a_{ij}}{\mu_{j}}+\frac{s_{ij}}{\mu_{j}}\right)^{2}$
$=$ $\sum_{j=1}^{m}\frac{\mu_{j}}{2}\left(\frac{s_{ij}}{\mu_{j}}\right)^{2}+\sum_{j=1}^{m}\frac{\mu_{j}}{2}\left(\frac{\lambda_{i}^{2}a_{ij}^{2}}{\mu_{j}^{2}}+\frac{2\lambda_{i}a_{ij}s_{ij}}{\mu_{j}^{2}}\right)$
$=$ $\sum_{j=1}^{m}\frac{\mu_{j}}{2}\left(\frac{s_{ij}}{\mu_{j}}\right)^{2}+\sum_{j=1}^{m}\lambda_{i}a_{ij}\left(\frac{\lambda_{i}a_{ij}}{2\mu_{j}}+\frac{s_{ij}}{\mu_{j}}\right)$
$=$ $\sum_{j=1}^{m}\frac{\mu_{j}}{2}\left(\frac{s_{ij}}{\mu_{j}}\right)^{2}+D_{i}(a_{i},a_{-i}),$
where the last equality is obtained by the expression of the cost function (2) and the definition of $s_{ij}$. Using an identical argument, we have $\Phi(a^{\prime}_{i},a_{-i})=\sum_{j=1}^{m}\frac{\mu_{j}}{2}\big{(}\frac{s_{ij}}{\mu_{j}}\big{)}^{2}+D_{i}(a^{\prime}_{i},a_{-i})$. Finally, by subtracting this relation from the former one, we get the desired relation (6).
Since the static game is an exact potential game, it admits a pure NE. Moreover, any sequence of best-response updates, in which a player updates its strategy and strictly reduces its cost, will also reduce the potential function. As a result, the best-response dynamics will converge to a local minimum of the potential function, which must be a pure NE.
Remark 3.3.
A more intuitive way to see why $\Phi(a)$ serves a potential function is to note that
$\frac{\partial\Phi(a)}{\partial a_{ij}}=\frac{\partial D_{i}(a)}{\partial a_{ij}}=\lambda_{i}\big{(}\frac{\lambda_{i}a_{ij}+s_{ij}}{\mu_{j}}\big{)}\ \forall i,j.$
This in view of [43, Lemma 4.4] suggests that the static load balancing game is a potential game.
Remark 3.4.
One can show that the potential function $\Phi(a)$ in the proof of Theorem 1 is strongly convex with respect to each player’s action separately. However, in general, $\Phi(a)$ may not be jointly convex with respect to the entire action profile $a$. Otherwise, we would have a unique pure NE.
Next, we proceed to characterize the best response of each player in a closed-form. To that end, let us define $\hat{\mu}_{ij}$ to be the relative available processing rate of server $j$ that is viewed by player $i$, i.e.,
$\hat{\mu}_{ij}=\frac{\mu_{j}}{s^{0}_{j}+\sum_{k\neq i}\lambda_{k}a_{kj}}.$
Then, we can rewrite the cost function (2) as
$D_{i}(a)=\sum_{j=1}^{m}\lambda_{i}a_{ij}\left(\frac{1}{\hat{\mu}_{ij}}+\frac{\lambda_{i}a_{ij}}{2\mu_{j}}\right).$ (7)
Given fixed strategies of the other players $a_{-i}$, a simple calculation shows that the entries of the Hessian matrix of the cost function (7) with respect to $a_{i}$ are given by
$\frac{\partial^{2}D_{i}(a)}{\partial a_{ij}\partial a_{i\ell}}=\begin{cases}\frac{\lambda_{i}^{2}}{\mu_{j}}\ &\mbox{if}\ \ell=j\\ 0\ &\mbox{if}\ \ell\neq j.\end{cases}$
Therefore, the Hessian of $D_{i}(a)$ with respect to $a_{i}$ is a diagonal matrix with strictly positive diagonal entries, and hence, it is a positive-definite matrix. This shows that the cost function of each player is strongly convex with respect to its own action.
Now suppose player $i$ plays its best-response action to the fixed actions of the other players (including the initial loads on the servers). As in [38], we can write a convex program whose solution gives us the best strategy for player $i$:
$\text{minimize}\quad D_{i}(a)$ (8)
[…]
$\ \quad\qquad a_{ij}\geq 0\quad\forall j\in[m].$ (10)
Since the objective function $D_{i}(a)$ is strongly convex with respect to the decision variable $a_{i}$, we can use KKT optimality conditions to solve the optimization problem (8) in a closed-form. The result is stated in the following lemma.
Lemma 3.5.
Given player $i$, assume the servers are sorted according to their available processing rates, i.e., $\hat{\mu}_{i1}\geq\hat{\mu}_{i2}\geq\cdots\geq\hat{\mu}_{im}$. Then, the optimal solution of the convex program (8) is given by
$a_{ij}=\begin{cases}\frac{\mu_{j}}{\lambda_{i}}\left(\frac{\lambda_{i}+\sum_{k=1}^{c_{i}-1}\frac{\mu_{k}}{\hat{\mu}_{ik}}}{\sum_{k=1}^{c_{i}-1}\mu_{k}}-\frac{1}{\hat{\mu}_{ij}}\right)&1\leq j<c_{i},\\[10.0pt] 0&c_{i}\leq j\leq m,\end{cases}$ (11)
where $c_{i}$ is the smallest integer index satisfying
$\hat{\mu}_{ic_{i}}\left(\lambda_{i}+\sum_{k=1}^{c_{i}-1}\frac{\mu_{k}}{\hat{\mu}_{ij}}\right)\leq\sum_{k=1}^{c_{i}-1}\mu_{k}.$
Proof 3.6.
Let us consider the Lagrangian function for player $i$, which is given by:
$L_{i}(a,\alpha,\eta)=D_{i}(a)-\alpha\left(\sum_{j=1}^{m}a_{ij}-1\right)-\sum_{j=1}^{m}\eta_{j}a_{ij},$
where $\alpha$ and $\eta=(\eta_{1},\ldots,\eta_{m})$ are Lagrange multiplies corresponding to the constraints in (8). Since (8) is a convex program, using KKT optimality conditions, we know that $a_{i}$ is the optimal solution (for a fixed $a_{-i}$) if and only if:
$\frac{\partial L_{i}}{\partial a_{ij}}=0\ \forall j,\quad\frac{\partial L_{i}}{\partial\alpha}=0,$
$a_{ij}\eta_{j}=0,\quad a_{ij}\geq 0,\quad\eta_{j}\geq 0\quad\forall j.$
Using the first condition, we have:
$\frac{\partial L_{i}}{\partial a_{ij}}=\frac{\lambda_{i}}{\hat{\mu}_{ij}}+\frac{\lambda_{i}^{2}a_{ij}}{\mu_{j}}-\alpha-\eta_{j}=0.$
If $a_{ij}=0$, we know that $\eta_{j}\geq 0$, which results in:
$\alpha\leq\frac{\lambda_{i}}{\hat{\mu}_{ij}}.$
If $a_{ij}\neq 0$, we have
$\eta_{j}=0\implies\alpha=\lambda_{i}\left(\frac{1}{\hat{\mu}_{ij}}+\frac{\lambda_{i}a_{ij}}{\mu_{j}}\right).$
Suppose we sort the servers according to their available processing rates such that
$\hat{\mu}_{i1}\geq\hat{\mu}_{i2}\geq\dots\geq\hat{\mu}_{im}.$
Now, a key observation is that a server with a higher available processing rate should naturally have a higher fraction of jobs assigned to it. Under the assumption that the servers are sorted by their available processing rates, the load fractions on the servers also should follow the same ordering, i.e.,
$a_{i1}\geq a_{i2}\geq\cdots\geq a_{im}.$
Consequently, there may be cases where slower servers are assigned no jobs. This implies the existence of an index $c_{i}\in[n]$ such that
$a_{ij}=0\quad\text{for }i=c_{i},\ldots,n.$
Therefore, suppose player $i$ only distributes its job among the servers $1\leq j<c_{i}$. We have
$\alpha\sum_{k=1}^{c_{i}-1}\mu_{k}=\lambda_{i}\sum_{k=1}^{c_{i}-1}\frac{\mu_{k}}{\hat{\mu}_{ik}}+\lambda_{i}^{2}$
$\implies\alpha=\frac{\lambda_{i}(\lambda_{i}+\sum_{k=1}^{c_{i}-1}\frac{\mu_{k}}{\hat{\mu}_{ik}})}{\sum_{k=1}^{c_{i}-1}\mu_{k}}.$
Finally, for $1\leq j<c_{i}$, we have
$a_{ij}=\frac{\mu_{j}}{\lambda_{i}}\left(\frac{\lambda_{i}+\sum_{k=1}^{c_{i}-1}\frac{\mu_{k}}{\hat{\mu}_{ik}}}{\sum_{k=1}^{c_{i}-1}\mu_{k}}-\frac{1}{\hat{\mu}_{ij}}\right),$
where $c_{i}\in[n]$ is the smallest index satisfying
$\frac{\lambda_{i}+\sum_{k=1}^{c_{i}-1}\frac{\mu_{k}}{\hat{\mu}_{ik}}}{\sum_{k=1}^{c_{i}-1}\mu_{k}}\leq\frac{1}{\hat{\mu}_{ic_{i}}}.$
In Theorem 1, we established that the best-response dynamics in the static game converge to a pure NE. In fact, using the closed-form characterization of the best responses in Lemma 11, we can prove a stronger result regarding the convergence time of the best-response dynamics to a pure NE in the static game. More specifically, we can show that under a mild assumption on the initial strategies, if each player updates their strategy exactly once using the best-response update rule (11), the game will reach a pure NE. This result is stated formally in the following theorem.
Theorem 3.7.
Suppose in the static load balancing game each player is initially distributing a non-zero portion of its job on each server, i.e., $a_{ij}^{0}\neq 0,\ \forall i,j$. Then, the action profile obtained when players update their actions exactly once (in any order) using the best-response rule in (11) is a pure NE. In other words, the best-response dynamics converge to a pure NE in $n$ iterations.
Proof 3.8.
Please see Appendix A.1 for the proof.
4 Main Theoretical Results for the Dynamic Load Balancing Game
In this section, we consider the dynamic load balancing game and present our main theoretical results. As we saw in Section 3, the static load balancing game admits a pure NE, and iterative best-response dynamics converge to one such NE point. Therefore, a natural question arises: will following the best-response dynamics in the dynamic game also lead to a NE? The main result of this section is to show that this is indeed the case, and to establish a bound on the number of iterations required for the dynamics to converge to a fully balanced distribution of loads across the servers.
Algorithm 1 Best-Response Dynamics for the Dynamic Load Balancing Game
Upon arrival of a new job at time $t$ to some player $i$ with job length $\lambda_{i}$:
  • Player $i$ observes the state $s^{t}$ at time $t$, computes the available processing rates $\hat{\mu}^{t}_{j}=\frac{\mu_{j}}{s_{j}^{t}},j\in[m]$, and sorts them in an increasing order, where without loss of generality, we may assume $\hat{\mu}^{t}_{1}\geq\cdots\geq\hat{\mu}^{t}_{m}.$
  • Player $i$ computes its best response at time $t$ by solving the convex program $\min\{D_{i}(a_{i},s^{t}):a_{i}\in A_{i}\}$, whose solution is given by
    $a^{t}_{ij}=\begin{cases}\frac{\mu_{j}}{\lambda_{i}}\left(\frac{\lambda_{i}+\sum_{k=1}^{c-1}s^{t}_{k}}{\sum_{k=1}^{c-1}\mu_{k}}-\frac{s^{t}_{j}}{\mu_{j}}\right)&1\leq j<c,\\[10.0pt] 0&c\leq j\leq m,\end{cases}$
    where $c$ is the smallest index satisfying:
    $\hat{\mu}^{t}_{c}\left(\lambda_{i}+\sum_{k=1}^{c-1}s_{k}\right)\leq\sum_{k=1}^{c-1}\mu_{k}.$
  • Player $i$ schedules $\lambda_{i}a^{t}_{ij}$ portion of its job on server $j\in[m]$, and the game proceeds to the next step $t+1$.
Algorithm 1 describes the best-response dynamics for the dynamic load balancing game. In this algorithm, at each time $t$, the player receiving a job distributes it among the servers according to its best-response strategy relative to the current state. The following lemma establishes a key property of the update rule in Algorithm 1.
Lemma 4.9.
Suppose each player follows Algorithm 1 in the dynamic load balancing game. Then, there exists a finite time $t^{\prime}=\max_{j\in[m]}\lceil s^{0}_{j}/\mu_{j}\rceil$ after which any player must distribute its job among all the servers, i.e., $a^{t}_{ij}>0\ \forall i,j,t\geq t^{\prime}$.
Proof 4.10.
Following Algorithm 1, suppose at an arbitrary time $t$ player $i$ distributes its job to the set of servers indexed by $\{1,\dots,c-1\}$, while sending nothing to the servers in $\{c,\dots,m\}$, i.e.,
$a^{t}_{ij}=\begin{cases}\frac{\mu_{j}}{\lambda_{i}}\left(\frac{\lambda_{i}+\sum_{k=1}^{c-1}s^{t}_{k}}{\sum_{k=1}^{c-1}\mu_{k}}-\frac{s^{t}_{j}}{\mu_{j}}\right)&1\leq j<c,\\[10.0pt] 0&c\leq j\leq m,\end{cases}$ (12)
where $c$ is the smallest index satisfying:
$\hat{\mu}^{t}_{c}\left(\lambda_{i}+\sum_{k=1}^{c-1}s_{k}\right)\leq\sum_{k=1}^{c-1}\mu_{k}.$ (13)
At time $t$, let us define
$C^{t}=\frac{\lambda_{i}+\sum_{k=1}^{c-1}s^{t}_{k}}{\sum_{k=1}^{c-1}\mu_{k}},$ (14)
which in view of (12) implies that
$\frac{\lambda_{i}a_{ij}^{t}+s_{j}^{t}}{\mu_{j}}=C^{t}\quad 1\leq j<c.$ (15)
Therefore, according to (4), the states of the servers evolve as
$\frac{s_{j}^{t+1}}{\mu_{j}}=\frac{(\lambda_{i}a_{ij}^{t}+s_{j}^{t}-\mu_{j})^{+}}{\mu_{j}}=(C^{t}-1)^{+}\quad 1\leq j<c,$ (16)
and
$\frac{s_{j}^{t+1}}{\mu_{j}}=\frac{(s_{j}^{t}-\mu_{j})^{+}}{\mu_{j}}\geq(C^{t}-1)^{+}\quad c\leq j\leq m,$ (17)
where the last inequality is obtained using the update rule of Algorithm 1 as for $c\leq j\leq m$, we have $\frac{s^{t}_{j}}{\mu_{j}}\geq C^{t}$.
Now, either $C^{t}\leq 1$, in which case by (16) all the servers in $\{1,\dots,c-1\}$ become empty at time $t+1$, or $C^{t}>1$, in which case by (16) and (17), the servers would still have some load remaining on them at time $t+1$. In either case, using (16) and (17), we have
$\frac{s_{1}^{t+1}}{\mu_{1}}=\dots=\frac{s_{c-1}^{t+1}}{\mu_{c-1}}=(C^{t}-1)^{+}\leq\frac{s_{c}^{t+1}}{\mu_{c}}\leq\dots\leq\frac{s_{m}^{t+1}}{\mu_{m}},$ (18)
or equivalently,
$\hat{\mu}^{t+1}_{m}\leq\cdots\leq\hat{\mu}^{t+1}_{c}\leq\hat{\mu}^{t+1}_{c-1}=\cdots=\hat{\mu}^{t+1}_{1}.$ (19)
This shows that once a player distributes its job among a set of servers, those servers will continue to receive jobs in future iterations simply because the order of those servers in terms of available processing rates remains on the top of the list. More precisely, let us denote the set of servers that receive a job at time $t$ by $S^{t}=\{1,2,\ldots,c-1\}$, and suppose another player (say player $i^{\prime}$) updates at time $t+1$ and wants to distribute its job among the servers. Then, according to Algorithm 1, player $i^{\prime}$ would send its job to all the servers in $S^{t}$ because they all have the highest order in the list (19), and possibly more servers examined in the same order as (19). Therefore, the set of servers receiving a job at time $t+1$ is a superset of those that receive a job at time $t$, and they form a nested sequence of sets over time, i.e.,
$S^{t}\subseteq S^{t+1}\subseteq S^{t+2}\subseteq\ldots.$ (20)
Next, we argue that the nested sequence (20) must grow until it eventually includes all the servers, i.e., $S^{t^{\prime}}=[m]$ for some $t^{\prime}$. In particular, using (20), we must have $S^{t}=[m]\ \forall t\geq t^{\prime}$, which means that all the servers must receive a job after time $t^{\prime}$, completing the proof.
To show that the nested sequence (20) must grow, we first note that if a server $j$ becomes empty at some time $t$, i.e., $s_{j}^{t}=0$, then $\hat{\mu}_{j}^{t}=\infty$. Thus, server $j$ must receive a job at the next time step $t+1$, and so by the nested property (20), it must receive a job in all subsequent iterations. Since $c\notin S^{t}$, it means that in all iterations $1,2,\ldots,t$, server $c$ did not receive a new job and was only processing its initial load $s^{0}_{c}$ at rate $\mu_{c}$. Otherwise, server $c$ must have previously received a job and thus would be in the set $S^{t}$ by the nested property (20). Therefore, the load of server $c$ at time $t$ would be $s_{c}^{0}-t\mu_{c}$. As a result, server $c$ does not belong to $S^{t}$ for at most $t\leq\lceil s^{0}_{c}/\mu_{c}\rceil$ iterations, after which its load becomes zero and according to the above discussion, it will receive a job in the next iteration. By repeating this argument for any choice of server $c$, one can see that after at most $t^{\prime}=\max_{j\in[m]}\lceil s^{0}_{j}/\mu_{j}\rceil$ iterations, we must have $S^{t^{\prime}}=[m]$.
We are now ready to state the main result of this section.
Theorem 4.11.
If all the players follow Algorithm 1, after
$t^{\prime\prime}=O\Big{(}\max_{j}\lceil\frac{s^{0}_{j}}{\mu_{j}}\rceil\frac{\sum_{j}\mu_{j}+\sum_{j}s_{j}^{0}}{\sum_{j}\mu_{j}-\lambda_{\max}}\Big{)}$
iterations, [^4] We refer to Appendix A.2 for an alternative bound on the number of iterations $t^{\prime\prime}$, which may be better for a certain range of parameters. the state of each server becomes zero (i.e., the system becomes load-balanced). In particular, the strategies of the players converge to the pure NE of the static load balancing game with zero initial loads.
Proof 4.12.
From Lemma 4.9, we know that once a player begins sending jobs to all servers, all subsequent players will continue to do so for the remainder of the time. Consider an arbitrary time $t\geq t^{\prime}$, and assume that at time $t$ player $i$ receives a job and plays its best-response according to Algorithm 1. Since all servers receive jobs at time $t\geq t^{\prime}$, we have
$a^{t}_{ij}=\frac{\mu_{j}}{\lambda_{i}}\left(\frac{\lambda_{i}+\sum_{k=1}^{m}s^{t}_{k}}{\sum_{k=1}^{m}\mu_{k}}-\frac{s^{t}_{j}}{\mu_{j}}\right),$
$(C^{t}-1)^{+}=\frac{s_{1}^{t+1}+\dots+s_{m}^{t+1}}{\mu_{1}+\dots+\mu_{m}}.$
Moreover, if player $i^{\prime}$ schedules its job at time $t+1$, then
$C^{t+1}=\frac{\lambda_{i^{\prime}}+\sum_{k=1}^{m}s^{t+1}_{k}}{\sum_{k=1}^{m}\mu_{k}}=\frac{\lambda_{i^{\prime}}}{\sum_{k=1}^{m}\mu_{k}}+(C^{t}-1)^{+}.$ (21)
There are two possible cases:
  • If $C^{t}>1$, then $\sum_{j=1}^{m}s_{j}^{t+1}\neq 0$. As in (16) where $c$ is replaced by $m$, we have
    $\frac{s_{j}^{t+1}}{\mu_{j}}=\frac{(\lambda_{i}a_{ij}^{t}+s_{j}^{t}-\mu_{j})^{+}}{\mu_{j}}=(C^{t}-1)^{+}\ \forall j\in[m].$ (22)
    Now because $C^{t}>1$, we have $(C^{t}-1)^{+}>0$, which implies $(\lambda_{i}a_{ij}^{t}+s_{j}^{t}-\mu_{j})^{+}>0$. Therefore, $(\lambda_{i}a_{ij}^{t}+s_{j}^{t}-\mu_{j})^{+}\!=\!\lambda_{i}a_{ij}^{t}+s_{j}^{t}-\mu_{j}\ \forall j\in[m]$. Using this relation in (22) and summing over $j$ we get
    $\sum_{j=1}^{m}s_{j}^{t+1}=\sum_{j=1}^{m}s_{j}^{t}-(\sum_{j=1}^{m}\mu_{j}-\lambda_{i}),$
    where we note that by the stability assumption, we have $0<\sum_{j=1}^{m}\mu_{j}-\lambda_{i}$.
  • If $C^{t}\leq 1$, then $\sum_{j=1}^{m}s_{j}^{t+1}=0$, and thus using (21), we have $C^{t+1}=\frac{\lambda_{i^{\prime}}}{\sum_{k=1}^{m}\mu_{k}}<1$. By repeating this argument inductively, one can see that $\sum_{j=1}^{m}s^{t+k}_{j}=0\ \forall k\geq 1$.
Using the above cases, one can see that if $\sum_{j=1}^{m}s^{t+1}_{j}\neq 0$ for some $t\geq t^{\prime}$, then the value of $\sum_{j=1}^{m}s^{t}_{j}$ decreases by at least $\sum_{j=1}^{m}\mu_{j}-\lambda_{\max}>0$ in the next time step, where $\lambda_{\max}=\max_{i}\lambda_{i}$. Therefore, after a finite number of iterations
$t^{\prime\prime}$ :-$\coloneq t^{\prime}+\frac{\sum_{j=1}^{m}s^{t^{\prime}}_{j}}{\sum_{j=1}^{m}\mu_{j}-\lambda_{\max}}:$
$\leq t^{\prime}+\frac{\sum_{j=1}^{m}s^{0}_{j}+t^{\prime}\lambda_{\max}}{\sum_{j=1}^{m}\mu_{j}-\lambda_{\max}}$ (23)
$=O\left(\max_{j}\lceil\frac{s^{0}_{j}}{\mu_{j}}\rceil\frac{\sum_{j}\mu_{j}+\sum_{j}s_{j}^{0}}{\sum_{j}\mu_{j}-\lambda_{\max}}\right),$ (24)
we must have $\sum_{j=1}^{m}s^{t+1}_{j}=0\ \forall t\geq t^{\prime\prime}$. In particular, for any time $t\geq t^{\prime\prime}$, the best-response action for the updating player $i$ takes a simple form of
$a^{t}_{ij}=\frac{\mu_{j}}{\mu_{1}+\dots+\mu_{m}}\ \ \ j\in[m],\forall t\geq t^{\prime\prime},$
that is, each server receives a fraction of the job proportional to its service rate and the load on each server would become $\frac{\lambda_{i}}{(\mu_{1}+\dots+\mu_{m})}$.
Finally, we note that since $\sum_{j=1}^{m}s^{t}_{j}=0$ for $t\geq t^{\prime\prime}$, the dynamic load balancing game reduces to the static load balancing game with zero initial loads, i.e., $s^{0}_{j}=0\ \forall j\in[m]$. Since by Theorem 1, the best-response strategy in the static game converges to a pure NE, we obtain that $\lim_{t\to\infty}a^{t}_{ij}=\frac{\mu_{j}}{\mu_{1}+\dots+\mu_{m}}$ for any $i\in[n]$ and $j\in[m]$, which is precisely the pure NE of the static game with zero initial loads.
(a) One player distributes a job at a time.
(b) All players distribute jobs simultaneously.
(c) Sequentially playing best-response dynamics.
Parameters for Setting 1: $\mu=[0.9,0.8,0.4,0.01]$, $\lambda=[0.5,0.5,0.3,0.7]$, $s^{0}=[10,10,1,0.5]$.
(d) One player distributes a job at a time.
(e) All players distribute jobs simultaneously.
(f) Sequentially playing best-response dynamics.
Parameters for Setting 2: $\mu=[1.4,1.2,1,0.5]$, $\lambda=[0.5,0.5,0.3,0.7,0.9,0.1,0.6,0.2]$, $s^{0}=[10,10,1,10]$.
(g) One player distributes a job at a time.
(h) All players distribute jobs simultaneously.
(i) Sequentially playing best-response dynamics.
Parameters for Setting 3: $\mu=[2,2]$, $\lambda=[0.5,0.5,0.3,0.7,0.9,0.1,0.6,0.2]$, $s^{0}=[10,50]$.
Figure 2: Comparison of different job distribution strategies across three settings. Each column corresponds to a specific strategy, and each row corresponds to a specific setting. Parameters for each row are included in their respective captions.
(a) $\mu=[1,1.5,2,0.6]$, $\lambda=[5,5,5,5]$, $s^{0}=[10,10,10,10]$
(b) $\mu=[0.3,0.8,0.2,0.01]$, $\lambda=[1,0.8,0.9,0.7]$, $s^{0}=[10,10,1,0.5]$
Figure 3: One player distributes its job at a time.
5 Numerical Experiments
In this section, we present numerical results to evaluate the performance of the best-response dynamics under various settings of the static game, dynamic game, sequential update rule, and simultaneous update rule. In our simulations, we consider three different choice of parameters for the initial state, players’ job lengths, and servers’ service rates, where the parameters for each setting are provided in the captions of the corresponding figures. For instance, in the first row of Fig. 2, we consider a settings with four servers with initial loads $s^{0}_{1}=10,s_{2}^{0}=10,s_{3}^{0}=1,s_{4}^{0}=0.5$ and service rates $\mu_{1}=0.9,\mu_{2}=0.8,\mu_{3}=0.4,\mu_{4}=0.01$, and a set of four players with job lengths $\lambda_{1}=0.5,\lambda_{2}=0.5,\lambda_{3}=0.3,\lambda_{4}=0.7$.
First, we simulate the performance of Algorithm 1 for the dynamic load balancing game, where at each time step, a single player is randomly selected to distribute its job across the servers. The results of this experiment are illustrated in Figs. 2(a), 2(d), 2(g), 3(a), and 3(b). As can be seen, after sufficient time has elapsed, all states converge to zero. Once the load on all servers becomes zero, it remains at zero, indicating a stable NE. As illustrated in these figures, the convergence of the dynamics to the zero state appears to be fast and scales polynomially with respect to the game parameters, which also justifies the theoretical performance guarantee provided by Theorem 4.11.
Next, we evaluate an alternative version of Algorithm 1 for the dynamic load balancing game, in which, at each time step, all players simultaneously distribute their jobs. In this case, each player adopts the best-response strategy, considering both the current state of the system and the strategies of other players. The results of this experiment are depicted in Figs. 2(b), 2(e), and 2(h). Interestingly, the results show that the simultaneous update of the players does not hurt the overall performance of the system and the state of the game eventually converges to zero, albeit at a slower convergence rate.
Finally, in Figs. 2(c), 2(f), and 2(i), we evaluate the performance of the sequential best-response dynamics in the static load balancing game, where $a^{0}_{ij}=\frac{1}{m}$ for all $i\in[n],j\in[m]$. As can be seen, the action profiles converge to a pure Nash equilibrium (NE) after exactly $n$ iterations (one update by each player), which confirms the theoretical convergence rate bound provided in Theorem 3.7.
6 Conclusions
In this work, we develop a game-theoretic framework for load balancing in static and dynamic settings, motivated by practical applications in job scheduling and cyber-physical systems. We demonstrate that the static game formulation is a potential game, which guarantees the existence of a pure Nash equilibrium, with best-response dynamics converging in a finite number of iterations. To account for the carryover effect, we extend our formulation to dynamic settings and show that players employing best-response strategies in each state can achieve a fully load-balanced system, with convergence to the Nash equilibrium of the static game occurring in polynomial time. Our numerical experiments validate the efficiency of the proposed algorithms.
These findings provide a solid foundation for further exploration into more complex load-balancing strategies in distributed systems. One promising direction for future work is to explore scenarios where all players update their strategies and distribute jobs simultaneously in each round of the dynamic game. While our numerical experiments suggest that the dynamic game would still converge to a load-balanced static game, the convergence rate is notably slower. Another interesting direction is to extend our work to stochastic settings, where servers have stochastic processing rates and process the jobs assigned to them with certain probabilities.
Appendix A Omitted Proofs
A.1 Proof of Theorem 3.7
Proof A.13.
Without loss of generality and by relabeling the players, let us assume that players update their strategies according to their index, i.e., the $i$th updating player is player $i$. Moreover, let us define $S^{t}_{i}$ to be the set of servers that receive a job from player $i$ given its action profile $a_{i}^{t}$ at time $t$. First, we show that if players sequentially update their actions according to their best responses, then $S^{1}_{1}\subseteq S^{2}_{2}\cdots\subseteq S^{n}_{n}$.
Consider an arbitrary player $i$ who wants to update its action at time $t=i$. Since players $\{i+1,\ldots,n\}$ have not updated their actions yet, player $i$ plays its best response with respect to the action profile $\{a_{1}^{1},\cdots,a_{i-1}^{i-1},a_{i+1}^{0},\cdots,a_{n}^{0}\}$. For each server $j$, define
$\hat{\mu}^{t}_{ij}=\frac{\mu_{j}}{\sum_{k=1}^{i-1}\lambda_{k}a^{k}_{kj}+\sum_{k=i+1}^{n}\lambda_{k}a^{0}_{kj}+s_{j}^{0}},$
which is the relative available processing rate of server $j$ that is viewed by player $i$ at time $t=i$. Thus, player $i$ sorts servers in a way that $\hat{\mu}^{t}_{i1}\geq\hat{\mu}^{t}_{i2}\cdots\geq\hat{\mu}^{t}_{im}$. Now suppose player $i$ sends a job to servers indexed in $\{1,\cdots,c_{i}-1\}$. We therefore have for $1\leq j<c_{i}$:
$C^{t}_{i}=\frac{\lambda_{i}a^{t}_{ij}}{\mu_{j}}+\frac{1}{\hat{\mu}^{t}_{ij}},$
where $c_{i}$ is the smallest integer index satisfying
$C_{i}^{t}\leq\frac{1}{\hat{\mu}^{t}_{i,c_{i}}}.$ (25)
Now, it is player $(i+1)$’s turn to update its policy and it calculates the available processing rate for each server. Thus, we have
$\frac{1}{\hat{\mu}^{t+1}_{i+1,j}}=\begin{cases}C_{i}^{t}-\frac{\lambda_{i+1}a^{0}_{i+1,j}}{\mu_{j}}&1\leq j<c_{i},\\[10.0pt] \frac{1}{\hat{\mu}^{t}_{ij}}-\frac{\lambda_{i+1}a^{0}_{i+1,j}}{\mu_{j}}&c_{i}\leq j\leq m.\end{cases}$ (26)
There are two possible scenarios:
  • $c_{i}-1=m$: This means that player $i$ distributes its job among all servers and we now want to check if player $i+1$ does the same thing. In this case we have
    $C_{i}^{t}=\frac{\lambda_{i+1}+\sum_{j=1}^{m}\frac{\mu_{j}}{\hat{\mu}^{t+1}_{i+1,j}}}{\sum_{j=1}^{m}\mu_{j}}.$
    Suppose server $k$ is the server with highest $\hat{\mu}^{t+1}_{i+1,k}$, so that player $i+1$ sends job to it. Now, we have to check whether player $i$ also sends a job to other servers. Take server $h\in[m]\setminus\{k\}$ as an example. Player $i+1$ sends a portion of its job to server $h$ only if
    $C_{i}^{t}<$ $\frac{\lambda_{i+1}a^{0}_{i+1,h}}{\mu_{h}}+\frac{\lambda_{i+1}+\frac{\mu_{k}}{\hat{\mu}^{t+1}_{i+1,k}}}{\mu_{k}}$
    $=$ $\frac{\lambda_{i+1}a^{0}_{i+1,h}}{\mu_{h}}+\frac{\lambda_{i+1}}{\mu_{k}}+C_{i}^{t}-\frac{\lambda_{i+1}a^{0}_{i+1,k}}{\mu_{k}},$ (27)
    which obviously holds because for all $i\in[n],j\in[m]$, $a_{ij}^{0}\neq 0$. More generally, for all $h\in[m]$ and $\mathcal{M}_{i+1}^{h}\subseteq[m]\setminus\{h\}$, we have
    $C_{i}^{t}<$ $\frac{\lambda_{i+1}a^{0}_{i+1,h}}{\mu_{h}}+\frac{\lambda_{i+1}+\sum_{j\in\mathcal{M}_{i+1}^{h}}\frac{\mu_{j}}{\hat{\mu}^{t+1}_{i+1,j}}}{\sum_{j\in\mathcal{M}_{i+1}^{h}}\mu_{j}}$
    $=$ $\frac{\lambda_{i+1}a^{t-1}_{i+1,h}}{\mu_{h}}+\frac{\lambda_{i+1}}{\sum_{j\in\mathcal{M}_{i+1}^{h}}\mu_{j}}+C_{i}^{t}$ (28)
    -- $\frac{\lambda_{i+1}\sum_{j\in\mathcal{M}_{i+1}^{h}}a^{0}_{i+1,j}}{\sum_{j\in\mathcal{M}_{i+1}^{h}}\mu_{j}}.$ (29)
    Thus, we conclude that player $c_{i+1}-1=m$, or in other words $S^{t}_{i}=S^{t+1}_{i+1}$.
  • $c_{i}-1<m$: Suppose player $i+1$ sends job to servers indexed in $\mathcal{N}^{t+1}_{i+1}\subseteq\{c_{i},\cdots,m\}$. Note that $\mathcal{N}^{t+1}_{i+1}$ can also be an empty set. We now want to see if it sends any job to servers indexed in $S^{t}_{i}$. Take server $k\in S_{i}^{t}$ as an example. Player $i+1$ sends job to server $k$ only if
    $C_{i}^{t}<$ $\frac{\lambda_{i+1}a^{0}_{i+1,k}}{\mu_{k}}+\frac{\lambda_{i+1}+\sum_{j\in\mathcal{N}^{t+1}_{i+1}}\frac{\mu_{j}}{\hat{\mu}^{t+1}_{i+1,j}}}{\sum_{j\in\mathcal{N}^{t+1}_{i+1}}\mu_{j}}$
    $=$ $\frac{\lambda_{i+1}a^{0}_{i+1,k}}{\mu_{k}}+\frac{\lambda_{i+1}-\lambda_{i+1}\sum_{j\in\mathcal{N}^{t+1}_{i+1}}a^{0}_{i+1,j}}{\sum_{j\in\mathcal{N}^{t+1}_{i+1}}\mu_{j}}$ (30)
    $+$ $\frac{\sum_{j\in\mathcal{N}^{t+1}_{i+1}}\frac{\mu_{j}}{\hat{\mu}^{t}_{ij}}}{\sum_{j\in\mathcal{N}^{t+1}_{i+1}}\mu_{j}}.$ (31)
    According to (25), we already know that
    $C_{i}^{t}\leq\frac{\sum_{j\in\mathcal{N}^{t+1}_{i+1}}\frac{\mu_{j}}{\hat{\mu}^{t}_{i,j}}}{\sum_{j\in\mathcal{N}^{t+1}_{i+1}}\mu_{j}}.$ (32)
    Since $a^{0}_{i+1,j}\neq 0$ for all $j\in[m]$, we conclude that
    $\frac{\lambda_{i+1}a^{0}_{i+1,k}}{\mu_{k}}+\frac{\lambda_{i+1}-\lambda_{i+1}\sum_{j\in\mathcal{N}^{t+1}_{i+1}}a^{0}_{i+1,j}}{\sum_{j\in\mathcal{N}^{t+1}_{i+1}}\mu_{j}}>0,$
    and therefore player $i+1$ sends a fraction of its job to server $k$ at time $t+1$. Now, to check whether player $i+1$ also sends a job to the rest of the servers in $S_{i}^{t}$, let us take an arbitrary server $h\in S_{i}^{t}\setminus\{k\}$. Then, player $i+1$ sends a job to server $h$ only if
    $C_{i}^{t}<$ $\frac{\lambda_{i+1}a^{0}_{i+1,h}}{\mu_{h}}+\frac{\lambda_{i+1}+\frac{\mu_{k}}{\hat{\mu}^{t+1}_{i+1,k}}+\sum_{j\in\mathcal{N}^{t+1}_{i+1}}\frac{\mu_{j}}{\hat{\mu}^{t+1}_{i+1,j}}}{\mu_{k}+\sum_{j\in\mathcal{N}^{t+1}_{i+1}}\mu_{j}}$
    $=$ $\frac{\lambda_{i+1}a^{0}_{i+1,h}}{\mu_{h}}$ (33)
    $+$ $\frac{\lambda_{i+1}+C_{i}^{t}\mu_{k}-\lambda_{i+1}a_{i+1,k}^{0}}{\mu_{k}+\sum_{j\in\mathcal{N}^{t+1}_{i+1}}\mu_{j}}$ (34)
    $+$ $\frac{\sum_{j\in\mathcal{N}^{t+1}_{i+1}}\left(\frac{\mu_{j}}{\hat{\mu}^{t}_{i,j}}-\lambda_{i+1}a^{0}_{i+1,j}\right)}{\mu_{k}+\sum_{j\in\mathcal{N}^{t+1}_{i+1}}\mu_{j}}.$ (35)
    In other words, we have to check whether
    $C_{i}^{t}<$ $\left(\frac{\mu_{k}+\sum_{j\in\mathcal{N}^{t+1}_{i+1}}\mu_{j}}{\sum_{j\in\mathcal{N}^{t+1}_{i+1}}\mu_{j}}\right)\left(\frac{\lambda_{i+1}a^{0}_{i+1,h}}{\mu_{h}}\right)$
    $+$ $\lambda_{i+1}\left(\frac{1-a_{i+1,k}^{0}-\sum_{j\in\mathcal{N}^{t+1}_{i+1}}a^{0}_{i+1,j}}{\sum_{j\in\mathcal{N}^{t+1}_{i+1}}\mu_{j}}\right)$ (36)
    $+$ $\frac{\sum_{j\in\mathcal{N}^{t+1}_{i+1}}\frac{\mu_{j}}{\hat{\mu}^{t}_{ij}}}{\sum_{j\in\mathcal{N}^{t+1}_{i+1}}\mu_{j}}.$ (37)
    More generally, for any $h\in S_{i}^{t}\setminus\{k\}$ and $\mathcal{H}_{i}^{h}\subseteq S_{i}^{t}\setminus\{h,k\}$, we have
    $C_{i}^{t}<$ $\left(\frac{\sum_{k\in\mathcal{H}_{i}^{h}}\mu_{k}+\sum_{j\in\mathcal{N}^{t+1}_{i+1}}\mu_{j}}{\sum_{j\in\mathcal{N}^{t+1}_{i+1}}\mu_{j}}\right)\left(\frac{\lambda_{i+1}a^{0}_{i+1,h}}{\mu_{h}}\right)$
    $+$ $\lambda_{i+1}\left(\frac{1-\sum_{k\in\mathcal{H}_{i}^{h}}a_{i+1,k}^{0}-\sum_{j\in\mathcal{N}^{t+1}_{i+1}}a^{0}_{i+1,j}}{\sum_{j\in\mathcal{N}^{t+1}_{i+1}}\mu_{j}}\right)$ (38)
    $+$ $\frac{\sum_{j\in\mathcal{N}^{t+1}_{i+1}}\frac{\mu_{j}}{\hat{\mu}^{t}_{ij}}}{\sum_{j\in\mathcal{N}^{t+1}_{i+1}}\mu_{j}},$ (39)
    which holds because of (32) and the fact that $a^{0}_{i+1,j}\neq 0$ for all $j\in[m]$. Since $i\in\{1,\cdots,n-1\}$ is chosen arbitrarily, we can conclude that $S^{1}_{1}\subseteq S^{2}_{2}\cdots\subseteq S^{n}_{n}$.
To complete the proof, we note that when player $n$ updates its action, it plays its best response with respect to the action profile $\{a_{1}^{1},a_{2}^{2},\cdots,a_{n-1}^{n-1}\}$. Suppose it sends a portion of its job only to servers indexed in $\{1,\cdots c_{n}-1\}$. We therefore know that $S_{i}^{i}\subseteq\{1,\cdots c_{n}-1\}$ for $i\in\{1,\cdots n-1\}$. Moreover, we know that player $n$ distributes its job in a way such that
$C^{n}_{n}=\frac{\sum_{i=1}^{n}\lambda_{i}a^{i}_{ij}+s_{j}^{0}}{\mu_{j}}\ \ \forall j\in\{1,\cdots m\}.$
In other words, player $n$ ensures that the normalized load is equal across all servers receiving jobs. Consequently, for all players, the servers to which they allocate jobs have identical normalized loads, which is exactly what players achieve when playing their best-response strategies. As a result, no player has an incentive to change their policy, as this distribution aligns with their best-response strategy. Thus, $\{a_{1}^{1},a_{2}^{2},\cdots,a_{n}^{n}\}$ forms a pure NE of the static game.
A.2 An Alternative Bound for the Convergence Rate of the Dynamic Game
In this appendix, we provide an alternative bound for the convergence rate of Algorithm 1 in the dynamic load-balancing game, which, for certain ranges of game parameters, could be better than the convergence rate given in Theorem 4.11.
As before, the main idea is to analyze how quickly $\sum_{j=1}^{m}s_{j}^{t}$ approaches zero. Define $\lambda_{\text{min}}=\min_{i}\lambda_{i}$ and $\lambda_{\text{max}}=\max_{i}\lambda_{i}$. Similarly, let $\mu_{\text{min}}=\min_{j}\mu_{j}$ and $\mu_{\text{max}}=\max_{j}\mu_{j}$. To lower-bound the amount of decrease in $\big{|}\sum_{j=1}^{m}s_{j}^{t+1}-\sum_{j=1}^{m}s_{j}^{t}\big{|}$, we consider two cases:
  • Not all servers are receiving jobs. In this case, the minimum change occurs when all servers, except the slowest one (the server with the lowest processing rate), receive and process jobs completely. For these servers, we have: $s_{j}^{t+1}-s_{j}^{t}=0$. Thus, only the slowest server contributes to the change in $\sum_{j=1}^{m}s_{j}^{t}$. Without loss of generality, suppose the index of this server is $m$, i.e., $\mu_{\text{min}}=\mu_{m}$. There are two possible cases:
    • (1)
      $\mu_{\text{min}}\leq s_{m}^{t}$: In this case we have $s_{m}^{t+1}=s_{m}^{t}-\mu_{\text{min}}$. This gives the inequality
      $\mu_{\text{min}}\leq\Big{|}\sum_{j=1}^{m}s_{j}^{t+1}-\sum_{j=1}^{m}s_{j}^{t}\Big{|}.$
    • (2)
      $s_{m}^{t}<\mu_{\text{min}}$: In this case we have $s_{m}^{t+1}=0$, thus $s_{m}^{t+1}-s_{m}^{t}=-s_{m}^{t}$. Moreover, since server $m$ has not received any job at time $t$, according to (13), we have $\frac{s_{m}^{t}}{\mu_{\text{min}}}\geq\frac{\lambda_{\text{min}}}{\sum_{j=1}^{m-1}\mu_{j}}$. This gives us
      $\frac{\mu_{\text{min}}\lambda_{\text{min}}}{\sum_{j=1}^{m}\mu_{j}-\mu_{\text{min}}}\leq\Big{|}\sum_{j=1}^{m}s_{j}^{t+1}-\sum_{j=1}^{m}s_{j}^{t}\Big{|}.$
  • All servers are receiving jobs, but the state has not yet converged to zero: According to Theorem 4.11, in this case we have
    $\sum_{j=1}^{m}s_{j}^{t}-\sum_{j=1}^{m}s_{j}^{t+1}=\sum_{j=1}^{m}\mu_{j}-\lambda_{i},$
    where $i$ is the player sending job at time $t$. Therefore, we have
    $\sum_{j=1}^{m}\mu_{j}-\lambda_{\text{max}}\leq\Big{|}\sum_{j=1}^{m}s_{j}^{t+1}-\sum_{j=1}^{m}s_{j}^{t}\Big{|}.$
Thus, putting all of the above cases together, we conclude that the number of iterations until convergence is upper-bounded by:
$t^{\prime\prime}$ $\leq\frac{\sum_{j=1}^{m}s_{j}^{0}}{|\sum_{j=1}^{m}s_{j}^{t+1}-\sum_{j=1}^{m}s_{j}^{t}|}$
$\leq\frac{\sum_{j=1}^{m}s_{j}^{0}}{\min\big{\{}\sum_{j=1}^{m}\mu_{j}-\lambda_{\max},\ \mu_{\min},\ \frac{\mu_{\text{min}}\lambda_{\min}}{\sum_{j=1}^{m}\mu_{j}-\mu_{\min}}\big{\}}}.$